# Emergent Symbolic Mechanisms Support

### v1

@@ 3.3 Representational Similarity Analyses
We also performed representational similarity analyses (Kriegeskorte et al., 2008) to better understand the representations that were produced by each type of attention head. In this analysis, representations are modeled in terms of their similarity with one another. For each set of tokens $A_{1},B_{1},...A_{N},A_{N}$, we created four prompts, intended to dissociate representations of abstract variables (i.e., symbols) vs. literal tokens. The first two prompts were the same as the $c_{1}^{abstract}$ and $c_{2}^{abstract}$ contexts used for causal mediation analysis, in which the same token plays different abstract roles. The other two prompts were based on $c_{1}^{abstract}$ and $c_{2}^{abstract}$, but the final instances of $A_{N}$ and $B_{N}$ were swapped. The resulting set of prompts predict one pattern of pairwise similarity for abstract variables, and a different pattern of similarity for literal tokens.
Figure LABEL:abstract_similarity shows the predicted pattern of pairwise similarities for representations of abstract variables, with all pairs involving two instances of the variable $A$ forming one block of high similarity, and all pairs involving two instances of $B$ forming another block. By contrast, Figure LABEL:token_similarity shows the predicted pattern of pairwise similarities for representations of literal tokens. This pattern displays 3 diagonal bands, corresponding to pairs of the same token (regardless of whether this token plays the same role).
@@
Interestingly, although symbol abstraction heads and symbolic induction heads primarily resemble the pattern predicted for abstract variables, they also show the diagonal bands predicted by the token similarity matrix, suggesting that they preserve some degree of token identity, and thus do not represent perfectly abstract variables. Nevertheless, the patterns of similarity displayed by these three distinct types of attention heads have a strikingly close correspondence to the predictions of the hypothesized architecture, with symbol abstraction heads and symbolic induction heads both generating outputs that primarily resemble abstract variables, and retrieval heads generating outputs that primarily resemble literal tokens.
@@ 3.4 Ablation Analyses
For each type of attention head, we performed a cumulative ablation analysis in which the heads with the top $h$ causal mediation scores were ablated. This was performed for the full range of $h=1...H$ heads in the entire model. We also performed a control experiment in which each head ablated in the previous experiment was replaced by the head in the same layer with the lowest causal mediation score. We measured the effects of these ablations in terms of the probability assigned to the correct answer.
We found that these ablation experiments had a dramatic effect for all three types of heads (Figures LABEL:abstraction_head_ablation-LABEL:retrieval_head_ablation). In the ablation condition, the probability assigned to the correct answer rapidly fell to zero as more heads were ablated, whereas in the control condition it was necessary to ablate almost all attention heads to have such an effect. These experiments confirmed that all three types of attention heads were both sufficient and necessary to perform the rule induction task.
@@ 3.5 Comparison with Induction Heads
For each attention head, we computed the prefix matching score previously used to identify induction heads (Olsson et al., 2022), and compared this with the causal mediation score for symbolic induction heads. We found that these scores were very weakly correlated ($r=0.13$), thus appearing largely orthogonal (Figure LABEL:induction_head_scatterplot). These results suggest that, despite the conceptual similarity between these two mechanisms, they are implemented by disjoint sets of attention heads.
@@ 3.6 Comparison with Function Vectors
We also investigated the relationship between symbolic induction heads and function vectors (Todd et al., 2023), representations of an in-context task that are generated by a subset of attention heads. The symbolic induction heads identified in the present work have many similarities to the attention heads that generate function vectors, including: 1) they are found in intermediate layers of the model, 2) they primarily attend to the last item in each in-context example, and 3) they are causally implicated in in-context learning for relational tasks.
@@
To address this, we computed the average indirect effect for each attention head, which represents a measure of the extent to which its outputs constitute function vectors (Todd et al., 2023), and compared this with the causal mediation score for symbolic induction heads. This analysis revealed that these scores are indeed highly correlated (Figure LABEL:function_vector_scatterplot; $r=0.82$), suggesting that they are essentially the same set of attention heads. That is, the output of symbolic induction heads can be thought of as function vectors. This result provides a novel perspective on function vectors, suggesting that, where relevant, they can be used to implement symbolic forms of computation. This result also provides insight into the mechanism that computes function vectors, suggesting that it may be conceptualized as a form of sequence induction over abstract variables.
@@ 4 Related Work
There is a rich history of work illustrating how various aspects of symbol processing might be implemented in neural networks. Work on the tensor product representation (Smolensky, 1990) and binding-by-synchrony (Hummel & Holyoak, 2003) illustrated how dynamic variable-binding can be performed in neural networks. Kriete et al. (2013) demonstrated how indirection, the use of one variable to refer to another, can be implemented in a biologically plausible neural network. More recently, a series of studies illustrated how a relational bottleneck (Webb et al., 2024b)–a strong inductive bias to perform relational processing–can enable data-efficient learning of abstract reasoning capabilities in deep learning systems (Webb et al., 2020; Kerg et al., 2022; Altabaa et al., 2023). The primary contribution of our work, relative to these previous studies, is to demonstrate empirically that symbolic mechanisms can emerge in a large-scale neural network, and to illustrate how they operate to support abstract reasoning. Notably, the symbol abstraction heads identified in this work implement an emergent version of the abstractor architecture that was previously proposed to support relational learning (Altabaa et al., 2023)
There has also been much recent work investigating the internal mechanisms that support various forms of abstract and structured task processing in language models. This work has identified key primitives such as induction heads (Olsson et al., 2022), function vectors (Todd et al., 2023), binding IDs (Feng & Steinhardt, 2023), and other mechanisms that play a role in relational processing (Merullo et al., 2023). We build on this previous work by identifying an integrated architecture that brings together multiple mechanisms. These include newly identified mechanisms – symbol abstraction and symbolic induction heads – that, respectively, carry out the processes of abstraction and rule induction needed to implement an emergent form of symbol processing that supports abstract reasoning in a neural network.
@@ 5 Discussion
On the one hand, the emergent architecture identified here, that supports abstract reasoning via an intermediate layer of symbol processing, is strikingly at odds with characterizations of language models as mere stochastic parrots (Bender et al., 2021) or ‘approximate retrieval’ engines (Wu et al., 2023). These results are also at odds with claims that neural networks will need innately configured symbol processing mechanisms in order to perform human-like abstract reasoning (Marcus, 2001; Dehaene et al., 2022; Wong et al., 2023). On the other hand, these results can be viewed as a vindication of longstanding claims that symbol-processing mechanisms of some form (whether they be innate or learned) are a necessary component supporting human cognitive abilities (Fodor & Pylyshyn, 1988), insofar as they suggest that neural networks can acquire these abilities by developing their own form of symbol processing.
@@
It is interesting to consider the extent to which the identified mechanisms are truly emergent vs. dependent on innate aspects of the model. The transformer architecture (Vaswani et al., 2017) does not obviously possess the strong relational inductive biases that characterize the abstractor (Altabaa et al., 2023) or other architectures designed to perform relational abstraction (Webb et al., 2024b). However, transformers do have some inductive biases that seem relevant, including: 1) an innate mechanism for computing in-context similarity via the inner product between keys and queries, and 2) a form of indirection, in the sense that the keys and queries that are used to select information for retrieval are distinct from the values that are retrieved. In future work, it would be interesting to investigate the extent to which these or other inductive biases contribute to the development of emergent symbol processing mechanisms.
@@
Finally, it should be noted that our results do not suggest the language model precisely implements the hypothesized architecture, but more likely that it implements an approximate version of these computations. For instance, the symbol abstraction head and symbolic induction head outputs do not appear to represent perfectly abstract variables, but rather contain some information about the specific tokens that are used in each problem. This might explain some of the ‘content effects’ that have been observed in language models (Dasgupta et al., 2022), in which reasoning performance is not entirely abstract, but depends on the specific content over which reasoning is performed. These effects are also well documented in human psychology (Wason, 1968), which suggests that a similar form of approximate symbol processing may be implemented by the human brain. Whether and how these systems can account for the reasoning mechanisms employed by the brain is an interesting and important question to explore in future work.
@@ Appendix A Code and Hardware
All code was written in Python using the TransformerLens and HuggingFace libraries. Experiments were conducted on two NVIDIA 80G H100 GPUs.
All code and data will be made publicly available upon acceptance.
@@ Appendix B Implementation Details
B.1 Task Performance Evaluation
We randomly selected English tokens from LLaMa-3 vocabulary to form 2000 prompts to evaluate the generation accuracy. The prompt format used in experiments (2-shot) is:
@@ B.2 Causal Mediation Analysis
1:  Input: context pair ($c_{1}$,$c_{2}$), language model $f(c)\in{R^{A}}$: outputs the logits for all possible next tokens at the last position of prompt $c$, vocabulary size $A$, the $i$-th value of vector $f(c)$: $f(c)[i]$, the correct answer for $c_{1}$: $y_{c_{1}}$, the expected answer for the patched context $c_{1}^{*}$: $y_{c_{1}^{*}}$.
2:  For $c_{1}$, measure the difference between the output logit for $y_{c_{1}^{*}}$ and $y_{c_{1}}$, i.e., $\Delta{f_{c_{1}}}=f(c_{1})[y_{c_{1}^{*}}]-f({c_{1}})[y_{c_{1}}].$
3:  Cache the inner activations after feeding $c_{2}$ into the model.
4:  Replace the activations of the entire attention block or individual attention head in certain layer with the activations from $c_{2}$ at specific token positions to intervene on processing of $c_{1}$, and retrieve the logit difference:
@@
The change in the logit difference is a measure of the causal mediation effect:
$s=\Delta{f_{c_{1}^{*}}}-\Delta{f_{c_{1}}}=(f(c_{1}^{*})[y_{c_{1}^{*}}]-f(c_{1}^{*})[y_{c_{1}}])-(f({c_{1}})[y_{c_{1}^{*}}]-f({c_{1}})[y_{c_{1}}]),$ (7)
5:  Output: Causal Mediation Score $s$
Algorithm 1 Causal Mediation
Algorithm 1 explains the causal mediation procedure on one context pair $(c_{1},c_{2})$. We randomly selected 20 prompts, restricting the analysis to prompts that the model answered correctly, and calculated the average score over both ABA and ABB tasks.
Contexts. For ($c_{1}^{abstract}$, $c_{2}^{abstract}$), $y_{c_{1}}=A_{N}$ (Eq. 1 and Eq.2). After activation patching, $y_{c_{1}^{*}}$ should be $B_{N}$ based on our hypothesis that the output embeddings of symbol abstraction heads and symbolic induction heads represent abstract variables. For ($c_{1}^{token}$, $c_{2}^{token}$), $y_{c_{1}^{*}}$ should be $B_{N}$ according to the hypothesis that the output of retrieval heads represent the actual token.
Activations. We focus on two types of activations, the output of the entire attention block and the output embeddings of the individual attention head. The attention block includes attention heads followed by an MLP. The output embedding of the attention heads is the average of the value embeddings weighted by the attention scores.
Positions. We replaced the activations at the final items of the in-context examples ($A_{n}$ or $B_{N}$) or the final token right before the generated tokens (the final separation token).

### v2

@@ 3.3 Representational Similarity Analyses
We also performed representational similarity analyses (Kriegeskorte et al., 2008) to better understand the representations that were produced by each type of attention head. In this analysis, representations are modeled in terms of their similarity with one another. For each set of tokens $A_{1},B_{1},...A_{N},B_{N}$, we created four prompts, intended to dissociate representations of abstract variables (i.e., symbols) vs. literal tokens. The first two prompts were the same as the $c_{1}^{abstract}$ and $c_{2}^{abstract}$ contexts used for causal mediation analysis, in which the same token plays different abstract roles. The other two prompts were based on $c_{1}^{abstract}$ and $c_{2}^{abstract}$, but the final instances of $A_{N}$ and $B_{N}$ were swapped. The resulting set of prompts predict one pattern of pairwise similarity for abstract variables, and a different pattern of similarity for literal tokens.
Figure LABEL:abstract_similarity shows the predicted pattern of pairwise similarities for representations of abstract variables, with all pairs involving two instances of the variable $A$ forming one block of high similarity, and all pairs involving two instances of $B$ forming another block. By contrast, Figure LABEL:token_similarity shows the predicted pattern of pairwise similarities for representations of literal tokens. This pattern displays 3 diagonal bands, corresponding to pairs of the same token (regardless of whether this token is assigned to the same variable).
@@
Interestingly, although the output embeddings for symbol abstraction heads and symbolic induction heads primarily resemble the pattern predicted for abstract variables, they also show the diagonal bands predicted by the token similarity matrix, suggesting that they preserve some degree of token identity, and thus do not exclusively represent abstract variables. However, it is still possible for abstract variables to be represented in an invariant manner in a subspace of these attention heads’ outputs. To test this, we performed a decoding analysis in which a decoder was trained to predict the abstract rule (ABA vs. ABB) for problems involving one set of tokens, and then tested on problems involving a completely different set of tokens. This decoding analysis yielded nearly perfect generalization accuracy for both symbol abstraction and symbolic induction heads ($>98\%$ test accuracy, see Section C.6 for more details), indeed suggesting that abstract variables are invariantly represented in a subspace of the output embeddings for these heads.
To gain a more precise understanding of the identified attention heads, we also applied RSA to the key, query, and value embeddings (Tables 2-3 and Figures LABEL:fig:_rsa_qkvo_abs-LABEL:fig:_rsa_qk_ind). For symbol abstraction heads, we found that queries primarily represented token identity, keys represented a mixture of both tokens and abstract variables, and values primarily represented the abstract variables. For symbolic induction heads, queries and keys primarily represented the relative position within each in-context example, while values primarily represented abstract variables. For retrieval heads, queries primarily represented abstract variables, keys represented a mixture of both tokens and variables, and values primarily represented the predicted token. These results further confirm the hypothesized mechanisms, namely that abstraction heads convert tokens to variables, symbolic induction heads make predictions over these variables, and retrieval heads convert variables back to tokens.
@@ 3.4 Ablation Analyses
For each type of attention head, we performed a cumulative ablation analysis in which the heads with the top $h$ causal mediation scores were ablated. This was performed for the full range of $h=1...H$ heads in the entire model. We performed a control experiment in which each head ablated in the previous experiment was replaced by the head in the same layer with the lowest causal mediation score. Additionally, we implemented a random baseline by ablating $h$ randomly selected attention heads. We measured the effects of these ablations in terms of the probability assigned to the correct answer.
We found that these ablation experiments had a dramatic effect for all three types of heads (Figures LABEL:abstraction_head_ablation-LABEL:retrieval_head_ablation). In the ablation condition, the probability assigned to the correct answer rapidly fell to zero as more heads were ablated, whereas in the control conditions it was necessary to ablate almost all attention heads to have such an effect. These experiments confirmed that all three types of attention heads were both sufficient and necessary to perform the rule induction task.
@@ 3.5 Comparison with Induction Heads
For each attention head, we computed the prefix matching score previously used to identify induction heads (Olsson et al., 2022), and compared this with the causal mediation score for symbolic induction heads. We found that these scores were very weakly correlated ($r=0.11$, Figure LABEL:induction_head_scatterplot). These results suggest that, despite the conceptual similarity between these two mechanisms, they are implemented by disjoint sets of attention heads.
@@ 3.6 Comparison with Function Vectors
We also investigated the relationship between symbolic induction heads and function vectors (Todd et al., 2024), representations of an in-context task that are generated by a subset of attention heads. The symbolic induction heads identified in the present work have many similarities to the attention heads that generate function vectors, including: 1) they are found in intermediate layers of the model, 2) they primarily attend to the last item in each in-context example, and 3) they are causally implicated in in-context learning for relational tasks.
@@
To address this, we computed the average indirect effect for each attention head, which represents a measure of the extent to which its outputs constitute function vectors (Todd et al., 2024), and compared this with the causal mediation score for symbolic induction heads. This analysis revealed that these scores are indeed highly correlated (Figure LABEL:function_vector_scatterplot; $r=0.86$), suggesting that they are essentially the same set of attention heads. That is, the output of symbolic induction heads can be thought of as function vectors. This result provides a novel perspective on function vectors, suggesting that, where relevant, they can be used to implement symbolic forms of computation. This result also provides insight into the mechanism that computes function vectors, suggesting that it may be conceptualized as a form of sequence induction over abstract variables. Interestingly, this is consistent with recent findings that function vector heads often develop out of induction heads (Yin & Steinhardt, 2025).
Additionally, we found that these heads are not the only mechanism involved in computing function vectors. Specifically, we found that symbol abstraction heads also play an important role. The previous analysis was based on function vector scores computed at the the final position in the sequence. When function vector scores were instead computed based on the third item in each context-example (Figure LABEL:fig:_func_cmp), these scores were highly correlated with symbol abstraction heads ($r=0.47$), but not symbolic induction heads ($r=0.06$). These results suggest that function vectors are first computed by symbol abstraction heads at the level of individual in-context examples, and symbolic induction heads are primarily responsible for aggregating them across in-context examples.
3.7 Evaluating Other Language Models
To determine how widespread the identified mechanisms are, we applied our causal mediation analyses to 12 additional open-source language models of different sizes across four model families: GPT-2, Gemma-2, Qwen2.5, and Llama 3.1 (Figure LABEL:model_summary). With the exception of the GPT-2 models, all tested LLMs displayed nearly perfect accuracy on the identity rules task with 10 in-context examples. Correspondingly, all of the LLMs that performed well on the identity rules task also displayed statistically significant scores for each type of attention head, and these heads were arranged in a three-stage hierarchy (Figure LABEL:fig:_model_head_num and Figures LABEL:fig:_three_heads_gemma2-LABEL:fig:_three_heads_llama). By contrast, the GPT-2 variants did not show robust evidence for the presence of symbol abstraction heads (Figure LABEL:fig:_model_head_num and LABEL:fig:_three_heads_gpt2), consistent with their poor performance on this task. These results further confirm the link between the identified mechanisms and abstract reasoning in LLMs, and suggest that symbolic mechanisms may only emerge at certain scales (in terms of model or training data size).
3.8 More Complex Reasoning Tasks
Our results thus far have focused on a simple but paradigmatic case of abstract reasoning – algebraic rule induction. Do these results also extend to more complex reasoning tasks? To test this, we extended our causal mediation analyses to two other abstract reasoning tasks (Figure 17). First, we investigated letter string analogies involving either successor or predecessor relations (Hofstadter et al., 1995). Second, we investigated verbal analogies involving either antonym or synonym relations. We found that the same basic three-stage architecture was present for both of these tasks (Figures LABEL:fig:_mech_lsa and  LABEL:fig:_mech_va), suggesting that similar mechanisms are indeed involved in more complex abstract reasoning tasks.
Interestingly, the causal mediation scores revealed a complex pattern of within- and between-task correlations. For the identity rules task and letter string analogies, we found moderate to high correlations both at the within-task level (i.e., for different relations or rules in the same task; Table 4) and the between-task level (Table 5). However, for verbal analogies, the correlations (both within- and between-task) were lower or even negative. These results suggest that, although all three tasks employed the same high-level mechanistic strategy, there was some specialization in the specific tasks and relations mediated by each attention head.
@@ 4 Related Work
There is a rich history of work illustrating how various aspects of symbol processing might be implemented in neural networks. Work on the tensor product representation (Smolensky, 1990) and binding-by-synchrony (Hummel & Holyoak, 2003) illustrated how dynamic variable-binding can be performed in neural networks. Kriete et al. (2013) demonstrated how indirection, the use of one variable to refer to another, can be implemented in a biologically plausible neural network. More recently, a series of studies illustrated how a relational bottleneck (Webb et al., 2024)–a strong inductive bias to perform relational processing–can enable data-efficient learning of abstract reasoning capabilities in deep learning systems (Webb et al., 2021; Kerg et al., 2022; Altabaa et al., 2024). The primary contribution of our work, relative to these previous studies, is to demonstrate empirically that symbolic mechanisms can emerge in a large-scale neural network, and to illustrate how they operate to support abstract reasoning. Notably, the symbol abstraction heads identified in this work implement an emergent version of the abstractor architecture that was previously proposed to support relational learning (Altabaa et al., 2024)
There has also been much recent work investigating the internal mechanisms that support various forms of abstract and structured task processing in large language models. This work has identified key primitives such as induction heads (Olsson et al., 2022), function vectors (Todd et al., 2024), concept vectors (Opiełka et al., 2025), binding IDs (Feng & Steinhardt, 2024), and other mechanisms that play a role in relational processing (Merullo et al., 2023). Complementary work has also investigated how structured mechanisms emerge in smaller models trained on synthetic tasks (Wu et al., 2025; Grant et al., 2025; Al-Saeedi & Härmä, 2025; Tang et al., 2025; Brinkmann et al., 2024). We build on this previous work by identifying an integrated architecture that brings together multiple mechanisms. These include newly identified mechanisms – symbol abstraction and symbolic induction heads – that, respectively, carry out the processes of abstraction and rule induction needed to implement an emergent form of symbol processing that supports abstract reasoning in neural networks.
@@ 5 Discussion
On the one hand, the emergent architecture identified here, that supports abstract reasoning via an intermediate layer of symbol processing, is strikingly at odds with characterizations of language models as mere stochastic parrots (Bender et al., 2021) or ‘approximate retrieval’ engines (Wu et al., 2024). These results are also at odds with claims that neural networks will need innately configured symbol processing mechanisms in order to perform human-like abstract reasoning (Marcus, 2001; Dehaene et al., 2022; Wong et al., 2023). On the other hand, these results can be viewed as a vindication of longstanding claims that symbol-processing mechanisms of some form (whether they be innate or learned) are a necessary component supporting human cognitive abilities (Fodor & Pylyshyn, 1988), insofar as they suggest that neural networks can acquire these abilities only by developing their own form of symbol processing.
@@
It is interesting to consider the extent to which the identified mechanisms are truly emergent vs. dependent on innate aspects of the model. The transformer architecture (Vaswani et al., 2017) does not obviously possess the strong relational inductive biases that characterize the abstractor (Altabaa et al., 2024) or other architectures designed to perform relational abstraction (Webb et al., 2024). However, transformers do have some inductive biases that seem relevant, including: 1) an innate mechanism for computing in-context similarity via the inner product between keys and queries, and 2) a form of indirection, in the sense that the keys and queries that are used to select information for retrieval are distinct from the values that are retrieved. In future work, it would be interesting to investigate the extent to which these or other inductive biases contribute to the development of emergent symbol processing mechanisms.
@@
Finally, an important open question concerns the extent to which language models precisely implement symbolic processes, as opposed to merely approximating these processes. In our representational analyses, we found that the identified mechanisms do not exclusively represent abstract variables, but rather contain some information about the specific tokens that are used in each problem. On the other hand, using decoding analyses, we found that these outputs contain a subspace in which variables are represented more abstractly. A related question concerns the extent to which human reasoners employ perfectly abstract vs. approximate symbolic representations. Psychological studies have extensively documented ‘content effects’, in which reasoning performance is not entirely abstract, but depends on the specific content over which reasoning is performed (Wason, 1968), and recent work has shown that language models display similar effects (Lampinen et al., 2024). In future work, it would be interesting to explore whether such effects are due to the use of approximate symbolic mechanisms, and whether similar mechanisms are employed by the human brain.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Code and Hardware
All code was written in Python using the TransformerLens and HuggingFace libraries. Experiments on Llama-3.1 70B and Qwen2.5 72B were conducted on two NVIDIA 80G H100 GPUs while experiments on other models of smaller sizes were conducted on a single H100 GPU. All model weights are loaded in the bfloat16 format.
All code and data will be released at https://github.com/yukang123/LLMSymbMech.
@@ Appendix B Implementation Details
B.1 Rule Induction Task: Evaluating Task Performance
To evaluate performance on the rule induction task, we randomly selected English tokens from the LLama-3 vocabulary to form 2,000 prompts. We used the following 2-shot prompt format:
@@ B.2 Causal Mediation Analysis
Algorithm 1 Causal Mediation Analysis
Input: context pair ($c_{1}$,$c_{2}$); language model $f(c)\in{R^{A}}$: outputs the logits for all possible next tokens at the last position of prompt $c$; vocabulary size $A$; the $i$-th value of vector $f(c)$: $f(c)[i]$; the correct answer for $c_{1}$: $y_{c_{1}}$; the expected answer for the patched context $c_{1}^{*}$: $y_{c_{1}^{*}}$.
1:  For $c_{1}$, measure the difference between the output logit for $y_{c_{1}^{*}}$ and $y_{c_{1}}$, i.e., $\Delta{f_{c_{1}}}=f(c_{1})[y_{c_{1}^{*}}]-f({c_{1}})[y_{c_{1}}].$
2:  Cache the internal activations after feeding $c_{2}$ into the model.
3:  Replace activations for selected model component (e.g., output of the entire attention block, or individual attention head, at a particular layer and token position) in $c_{1}$ with corresponding activations in $c_{2}$, yielding patched context $f_{c_{1}^{*}}$.
4:  Compute logit difference for patched context:
@@
5:  Compute causal mediation score:
$s=\Delta{f_{c_{1}^{*}}}-\Delta{f_{c_{1}}}=(f(c_{1}^{*})[y_{c_{1}^{*}}]-f(c_{1}^{*})[y_{c_{1}}])-(f({c_{1}})[y_{c_{1}^{*}}]-f({c_{1}})[y_{c_{1}}])$ (7)
Output: Causal Mediation Score $s$
Algorithm 1 explains the causal mediation procedure for a given context pair $(c_{1},c_{2})$. We performed three separate versions of this analysis, each targeting one of the hypothesized attention head mechanisms, as described below.
Targeting Symbol Abstraction Heads To target the hypothesized symbol abstraction heads, we used the context pair ($c_{1}^{abstract}$, $c_{2}^{abstract}$). In one version of this analysis, $c_{1}^{abstract}$ involved an ABA rule, and $c_{2}^{abstract}$ involved an ABB rule. Importantly, the same tokens were used for both contexts, but with swapped variable assignments, such that the correct answer to complete both contexts was the same. We also performed a version of this analysis in which $c_{1}^{abstract}$ was an ABB rule and $c_{2}^{abstract}$ was an ABA rule. We used 100 prompts for each version of the analysis, yielding 200 prompts in total. Patching was performed at the position of the final token in each of the two in-context examples. See Figure 6 for an illustration and further explanation of this analysis.
Targeting Symbolic Induction Heads. To target the hypothesized symbolic induction heads, we performed the same analysis described above (for symbol abstraction heads), except that patching was performed at the final position in the sequence, i.e. the position at which the model was expected to generate the missing token. See Figure 7 for an illustration of this analysis.
Targeting Retrieval Heads. To target retrieval heads, we used the context pair ($c_{1}^{token}$, $c_{2}^{token}$). In one version of this analysis, $c_{1}^{token}$ involved an ABA rule, and $c_{2}^{token}$ involved the same ABA prompt, but with the position of the final two tokens swapped. We also performed a version of this analysis in which both $c_{1}^{token}$ and $c_{2}^{token}$ involved an ABB rule. Patching was performed at the final position in the sequence. See Figure 8 for an illustration of this analysis.
Activations. We performed CMA on two types of activations: 1) the output of the entire attention block (including both the attention head and MLP outputs) at a specific layer and token position, and 2) the output of a specific attention head (before the additional projection matrix over concatenated head outputs) at a specific layer and token position. The first analysis was intended to provide a course-grained, but comprehensive, account of where the causally relevant computations were located (i.e., which layers and token positions). The second analysis was intended to provide a more fine-grained analysis (focusing on individual attention heads) of the causally relevant model components at those locations.
Permutation Testing. We performed statistical tests to identify significant heads for each head type. Specifically, we used permutation testing to model the null distribution for these analyses. For each attention head $h$, the average causal score over multiple context pairs measures the difference between the means of two groups, i.e., the logit difference for the original context $\Delta{f_{c_{1}}}$ and the logit difference for the patched context $\Delta{f_{c_{1}^{*}}}$. To model the null distribution, we randomly permuted these two quantities, such that there was a 50% chance they would remain the same, and a 50% chance $\Delta{f_{c_{1}}}$ would be swapped with $\Delta{f_{c_{1}^{*}}}$ and vice versa. This was repeated 5,000 times. We then identified the threshold $\epsilon$ for which there was a family-wise error rate of $p<0.05$. That is, we identified the threshold for which there was a less than 5% chance that even a single attention head would have a causal mediation score exceeding this threshold in the null distribution. We considered only heads with scores that passed this threshold to be statistically significant.
Figure 6: Illustration of Causal Mediation Analysis targeting Symbol Abstraction Heads. In this example, context $c_{1}$ involves an ABA rule, and context $c_{2}$ involves an ABB rule. Importantly, both contexts are formed from the same token sets, but with the tokens assigned to different variables, such that the correct prediction for the next token is the same in both contexts (‘ks’). Patching from context $c_{2}$ to $c_{1}$ should therefore not have an effect on model components that encode literal tokens (i.e., the retrieval heads), since the output of these components should be the same in both contexts. This intervention should, however, have an effect on model components that encode abstract variables, since these differ between the two contexts. To target symbol abstraction heads, we patched from $c_{2}$ to $c_{1}$ at the position of the final tokens in each in-context example, thus intervening on the abstract variable representation for these tokens (and therefore converting the representation of an ABA rule into a representation of an ABB rule). In this example, the correct answer $y_{c_{1}}$ for context $c_{1}$ is ‘ks’ (i.e., $A_{N}$), and the expected answer $y_{c_{1}^{*}}$ for the patched context $c_{1}^{*}$ is ‘ixe’ (i.e., $B_{N}$).
Figure 7: Illustration of Causal Mediation Analysis targeting Symbolic Induction Heads. The intervention targeting symbolic induction heads was the same as the intervention targeting symbol abstraction heads, except that patching was performed at the final position in the sequence, at which the model was expected to generate the final token. This intervention should have an effect on the process of predicting the abstract variable associated with the final token, but not on the process of retrieving the final token itself.
Figure 8: Illustration of Causal Mediation Analysis targeting Retrieal Heads. In this example, both contexts $c_{1}$ and $c_{2}$ involve an ABA rule, but the order of the tokens in the final in-context examples are swapped. This means that the correct answer differs between these two contexts, and therefore model components that encode literal tokens such as retrieval heads should be affected by this intervention. Model components that encode abstract variables, however, should not be affected by this intervention, since the abstract rule in both contexts is the same. To target Retrieval Heads, we patched from $c_{2}$ to $c_{1}$ at the final position in the sequence. In this example, the correct answer $y_{c_{1}}$ for context $c_{1}$ is ‘ks’ (i.e., $A_{N}$), and the expected answer $y_{c_{1}^{*}}$ for the patched context $c_{1}^{*}$ is ‘ixe’ (i.e., $B_{N}$).